OpenStudy (anonymous):

Find the complex zeros of the polynomial function. Write f in factored form f(x)=x^3+1 Reduce fractions and simplify roots

OpenStudy (tkhunny):

I'd try the factor (x+1) first.

OpenStudy (anonymous):

I have no clue how to do any part of this problem

OpenStudy (tkhunny):

Synthetic Division? Long Division? You need to have a clue unless you have been sleeping through class.

OpenStudy (anonymous):

I have looked at the examples given to me for the last 5 problems and I just get real confused

OpenStudy (anonymous):

tkhunny my class in all online

OpenStudy (jdoe0001):

hmmm \(x^3+1 \implies x^3+1^3\)

OpenStudy (tkhunny):

I can buy the online argument. That can be tricky. Did you study factoring?

OpenStudy (anonymous):

no not really. The class is only about 5wks long. it's done in 9 days

OpenStudy (jdoe0001):

I don't think online classes including not covering material before exercises

OpenStudy (campbell_st):

this question is the sum of 2 cubes... and the basic structure is \[a^3 + b^3 = (a +b)(a^2 -ab + b^2)\] you don't need polynomial or sythetic division just recognise a = x and b = 1 then substitute for the quadratic you will need to use the general quadratic formula to find the complex roots

OpenStudy (anonymous):

this is intro to college math 1

OpenStudy (anonymous):



OpenStudy (anonymous):

that is what the answer for the example looks like

OpenStudy (anonymous):

You have to solve this equation: x^3+1=0 or x^3=-1 One obvious solution is x=-1, since (-1)(-1)(-1)=-1 By división bt (x+1) you can find the rest : (x^2-x+1) so: x^3+1=(x+1)(x^2-x+1) to find the other 2 complex roots just use standart method of disciminant

OpenStudy (anonymous):

so type x^3+1=(x+1)(x^2-x+1) that into my website I use

OpenStudy (anonymous):

no. You are asked for the roots. Find the roots of x^2-x+1=0 now

OpenStudy (anonymous):

\(x=\huge \frac{1 \pm \sqrt{1-4}}{2}=\frac{1\pm i\sqrt{3}}{2}\)

OpenStudy (anonymous):

what is the + with the line under it

OpenStudy (anonymous):

plus minus

OpenStudy (anonymous):



OpenStudy (anonymous):

yea, i've seen it just didn't know what it meant

OpenStudy (anonymous):

so now the equation thing you gave is that a step closer to the answer?

OpenStudy (anonymous):

I gave you the answer. Complex roots are: (1/2)+i(sqrt3)/2 and (1/2)-i(sqrt3)/2

OpenStudy (anonymous):

the thing a bit above is wrong

OpenStudy (anonymous):

i don't think so

OpenStudy (anonymous):

\(x^3+1=(x+1)( \frac{1}{2}+i\frac{\sqrt3}{2})( \frac{1}{2}-i\frac{\sqrt3}{2})\)

OpenStudy (anonymous):

this is factored form that you are asked for